OpenStudy (anonymous):

find cos θ if sin θ =-5/13 and tan θ>0

OpenStudy (solomonzelman):

Okay. \[Tip~~1:~~~~~~~~~~5^2+12^2=13^2\] sin= negative, so it's either 2nd or 4th quadrant tan= opp/adj and it's more than zero.

OpenStudy (solomonzelman):

 which one, in the fourth or second quadrant or both? explain?